OpenStudy (anonymous):

2y = x + 3 5y = x - 7 What is the solution set of the given system? {(-1/3, -4/3)} {(1/3, 4/3)} {(-29/3, -10/3)} {(29/3, 10/3)}

OpenStudy (anonymous):

5x + 2y = 7---(1) y = x + 1---(2) Substitute y from (2) into (1) You get 5x + 2(x+1) = 7 5x + 2x + 2 = 7 7x = 5 x = 5/7 Substituting the value of x in (2) y = 5/7 + 1 = 12/7

OpenStudy (anonymous):

2y+-x=3 5y-x=-7 So the third option is correct. You can use elimination, substitution, or the amtrix on your calculator to solve. {(-29/3, -10/3)}

OpenStudy (anonymous):

2x - y - 4 = 0 3x + y - 9 = 0 What is the solution set of the given system? {(6/5, 13/5)} {(13/5, 6/5)} {(5, 6)} {(6, 5)}

OpenStudy (anonymous):

{(13/5, 6/5)}

OpenStudy (anonymous):

What is the result of adding the system of equations? 2x + y = 4 3x - y = 6 x= 2 x= 10 5x= 10